3.Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. Please show work
Answer by ewatrrr(24785)

 
Hi, Previously Posted
Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 ( m = -1)
New Line: Parallel..same slope
y = -x + b
2 = -(-4) + b
-2 = b
y = -x -2